# Scaling of poroelastic coarsening and elastic arrest in crosslinked gels

### v1

@@ Abstract
Recent experiments on crosslinked gels quenched from solvent-rich to solvent-poor conditions, show solvent-rich domains surrounded by gel-rich regions that coarsen followed by kinetic arrest at micron scales that persists for hours before full drainage to macroscopic equilibrium occurs on day timescales. Motivated by this, we present a general model for both coarsening and eventual arrest of pressure-driven coarsening in gels with solvent flow driven by interfacial tension of the solvent-rich domains and the gel. In gels in their viscoelastic time regime, this capillary force is dissipated by solvent transport through crosslinked polymers where the polymers develop transient elastic stresses due to solvent flow. At longer times, the long-ranged, elastic response of the gel provides a deterministic force which balances the capillary force thus arresting pressure-driven coarsening at a stiffness-dependent length. For a melt-like, polymer-rich gel, the coarsening length $\sim t^{1/4}$ where $t$ is the time, with an amplitude $\sim G^{-1/2}$, where $G$ is the shear modulus. The arrest length in this regime also scales $\sim G^{-1/2}$. For low polymer fractions where the dominant length scale is the mesh size, the coarsening length can scale as $t^{1/3}$ varying with $G^{-1/3}$ (as does the arrest length). Our predictions for the arrest-length scaling with $\sim G^{-1/2}$ for the melt-like gel are consistent with the measurements.
@@ I Introduction
Recent experiments [9] on phase separation of PDMS cross-linked polymers swollen by an organic solvent (a polymer gel system) have shown results that are strikingly different from our simple understanding of phase separation in simpler, two-component systems. The cross-linked polymer PDMS gel is initially prepared at a higher temperature with a certain fraction of solvent in equilibrium. The temperature is then lowered, and the equilibrium solvent fraction within the gel decreases. Once the gel comes to equilibrium—which only occurs macroscopically after a day or two—the gel and the solvent exhibit phase separation.
At its lower temperature state, the gel initially contains excess solvent due to its initial high-temperature preparation. After an initial period of spinodal decomposition and nucleation, mesoscale domains of solvent-rich regions form in the gel, surrounded by gel that is polymer-rich and has less solvent than its initial, high temperature state. The domain structure coarsens (with small domains shrinking and large ones growing), and this process is reminiscent of Lifshitz-Slyozov-Wagner (LSW) [15, 27, 3], coarsening in the phase separation of two-component systems. Strikingly, in these experiments the coarsening is observed to undergo long-lived kinetic arrest at submicron or micron length scales, depending on the rigidity of the cross-linked gel. This mesoscale structure remains stable for several hours until, on the scale of days, all excess solvent eventually drains from the solvent-rich domains into a reservoir. The gel then becomes homogeneous at its low-temperature state with a lower solvent fraction than its prepared state.
@@
A paper with theory and model calculations of phase separation and coarsening in a viscoelastic gel by Curk and Luijten [5] extended LSW-type Ostwald ripening driven by capillarity (interfacial tension $\sigma$) by incorporating elastic stresses and their viscoelastic relaxation in the surrounding matrix. Their approach is based on diffusion-controlled evaporation–condensation through a viscoelastic solid, modified by elastic and viscoelastic effects. They do not include poroelastic (Darcy) solvent flow through a deformable, porous network, which provides the dominant dissipative mechanism for relaxing capillary pressure in our approach and leads to the elastic arrest prediction of scaling with $1/G^{1/2}$, in agreement with the experiments [9]. A previous, numerical calculation by Onuki and Puri [17] focused on a two-dimensional crosslinked gel with (implicit) solvent-gel friction. In their numerical solutions the onset of elastic arrest occurred at a time that depended on the value of the modulus (see Appendix for further discussion). A more recent theory of phase separation in elastic networks [20], based on coupled chemical-potential gradients and elastic stress, concluded that local elasticity alone cannot arrest coarsening, and that non-local elastic interactions are required. By contrast, our approach focuses explicitly on solvent flow through the gel coupled to its poroelastic response. By treating solvent permeation (Darcy flow) and elastic back-reaction directly—rather than incorporating elasticity solely through constitutive stress–energy relations—we demonstrate that elastic arrest of coarsening can occur even for linearly elastic gels, as explained below.
@@
Here, we present a unified scaling theory of both domain coarsening and elastic arrest in crosslinked gels which exhibit crosslink-controlled elastic behavior at long times and a viscoelastic response activated by solvent flow through the network at intermediate times. Darcy flow [26, 2, 13] of the solvent in the gel occurs at time scales for which poroelastic dissipation, acting within the viscoelastic regime of the polymer network, controls the rate of coarsening, while the onset of long-time elastic response governs kinetic arrest [8, 17, 26]. When the characteristic coarsening time becomes comparable to the relaxation time separating flowing, viscoelastic regime from the solid-like elastic regime, further capillary-driven domain evolution is arrested by the development of long-range elastic stress in the gel. At the level of the Laplace force balance at the solvent-rich domain boundary, this elastic stress cancels the capillary driving force and halts further solvent flow. The resulting arrested domain size is therefore not set directly by microscopic gel length scales, but by a kinetic criterion and is amplified by a magnification factor arising from the large separation between the poroelastic time governing solvent transport and the much longer time required for the onset of elastic response. During this extended viscoelastic window, solvent can be transferred over progressively larger distances, converting nanoscale gel properties into mesoscale domain sizes at arrest.
The complexity of the problem is simplified by our scaling approach focused on the dimensionally correct dependence of the coarsening and elastic arrest in terms of the gel properties, without the fine distinctions of numerical factors or dimensionless factors of order unity [6]. Thus, instead of equalities, we use the convention $\sim$ to mean “scales as”. The theory predicts a coarsening exponent of $1/4$ for slow, poroelastic diffusion and the arrest size of the domains scales as $1/\sqrt{G}$; the latter agrees with the experiments. Finally, we note here that while the explanations below consider the case of shrinking, solvent-rich domains, our model and its scaling predictions apply to both shrinking and growing domains, which are of course related by conservation of solvent. When we discuss solvent flowing out of shrinking solvent-rich domains, the same physics applies to solvent being drawn into growing domains.
@@ II.1 LSW coarsening
These domains coarsen [3] and the kinetics of coarsening begins with the conservation relation that equates the number of molecules $N_{B}$ leaving (for example) a shrinking B domain (and eventually being drawn into a growing B domain), to their flux $J_{BA}$ at the interface of the B domain with the background of A: $dN_{B}/dt=-J_{BA}S$, where the flux is evaluated locally at the domain boundary and is controlled by the driving force acting at that boundary. For scaling purposes, we can ignore the vectors, keeping in mind that the flux and velocity of the B-molecules exiting the B-rich domain are in the direction normal to the domain boundary. Here $S$ is the area of the shrinking domain which, for simplicity, we have taken to be spherical and henceforth we denote the time derivatives by a dot. $\dot{N}_{B}$ is related to the composition difference $\Delta c$ between the two phases and the change in the domain interface position and the domain area. The flux is the number of B molecules that exit the B-rich domain, per unit area per unit time and is written in terms of the B-molecule velocity as: $J_{BA}=c_{B}\ v$, where $c_{B}$ is the concentration (number density) of B molecules in the B-rich domain. For a shrinking domain, $\dot{N}_{B}<0$ implies that $J_{BA}>0$; the velocity is directed away from the domain. In scaling, we write: $\dot{N}_{B}\sim-\Delta c\,S\,\dot{\lambda}$ where $\lambda$ is the size of the $B$ domain. For a more detailed discussion of this and the following see Appendix C.
@@
Using the conservation equation, we now write a scaling relation: $\Delta c\ \dot{\lambda}\sim-D\ \nabla\sigma/\lambda$ where $\Delta c$ is the concentration difference of the B molecules in the B-rich phase and the A-rich phase. $\nabla$ now is the extent of the chemical potential gradient in the direction normal to the surface of the B domain, evaluated at the domain interface. LSW scaling for the domain growth is obtained by approximating the gradient term by $\sigma/\lambda^{2}$ which originates in the solution of the diffusion equation for the B molecules in the background of A, at the domain surface [3]. This equation in the fast-diffusion limit (applicable to molecular diffusion in a liquid relative to the coarsening times), results in an equation for the spatial profile of the B molecule concentration in the A-rich domain where $\lambda$ determines the scale of the gradient. Below, we generalize the approximation for the gradient to include diffusion.
@@
The result for a shrinking domain gives $\dot{\lambda}\sim-(D/\Delta c)\,\sigma/\lambda^{2}$. For scaling, we can take $\Delta c\sim c_{B}$ (not valid near a critical point) which gives the LSW scaling, $\lambda\sim\left(\ (D\ \sigma/c_{B})\ t\right)^{1/3}$, where we do not distinguish shrinking from growing vis a vis the scaling; the two are simply related by conservation of the B molecules. As discussed below, the equations for the chemical potential gradients can be transformed into equations for the osmotic pressure gradients of the B molecules in the A-rich phase.
II.2 Generalizing LSW to coarsening of solvent-rich domains in gels:
@@
In the following, the gel properties such as interfacial tension with a solvent-rich domain, elastic modulus $G$, mesh size $\xi$ are envisioned as those that characterize the equilibrium gel. We again begin with the conservation relation, $\dot{N}_{B}\sim-\Delta c_{s}\,S\,\dot{\lambda}=-JS$ where the flux is evaluated locally at the domain boundary and is controlled by the driving force acting at that boundary. $\Delta c_{s}$ is the difference between the solvent concentration in the solvent-rich domain and the gel. The flux is given by $\vec{J}=c_{s}\vec{v}$ where $c_{s}$ solvent concentration in the solvent-rich domain which, due to capillary pressure, ejects solvent molecules into the gel with a $\vec{v}$. For a more detailed discussion of this and the following see Appendix C.
@@
However, instead of diffusing in the polymer-rich phase, the velocity is governed by solvent flow in the deformable gel considered as a porous medium where the Darcy relation [6, 1, 2, 24] that connects the velocity to the pressure gradient: $\vec{v}=(k_{0}/\eta_{D})\nabla p$ where $k_{0}$ is the permeability of the gel and $\eta_{D}$ is the viscosity appropriate to the solvent flow impeded by its friction with the polymer network. Because this transport is controlled by the hydrodynamic permeability of the solvent rather than by polymer rearrangements, the network topology is effectively fixed on the Darcy-flow timescale: polymer configurations do not relax rapidly enough to eliminate the stress induced by solvent flow. Physically, this reflects a separation of timescales between the relatively fast permeation of small solvent molecules through the gel and the much slower relaxation of stress carried by long-chain polymers. In the Results, we discuss specific choices for these parameters.
@@
The pressure gradient entering Darcy’s law is the local gradient at the interface of the solvent-rich domain and the gel, whose magnitude is set by how the interfacial pressure perturbation spreads into the gel. In incompressible systems pressure gradients are proportional to chemical potential gradients of the solvent molecules in the gel, by the thermodynamic relation, $\nabla\mu=\nabla p/c_{B}$; see Eq. 1.1.14 of Ref. [19] and Appendix C. The chemical potential at the solvent-rich domain surface is again given by capillarity, $\delta\mu=\sigma/(\lambda c_{s})$ and the osmotic pressure jump at the interface (which acts as the mechanical pressure jump) is given by the Laplace law as $\delta p=c_{s}\delta\mu=\sigma/\lambda$. Finally, the conservation law and the Darcy relation used in the flux give the dynamical scaling relation for a shrinking domain:
$\dot{\lambda}\sim-\frac{k_{0}\ c_{s}}{\eta_{D}}\nabla p\sim-\frac{k_{0}\ \sigma}{\eta_{D}\ \lambda\ \ell_{p}}$ (1)
@@
The solvent pressure in the gel decreases from its maximal value at the boundary of the shrinking domain via stress diffusion, known as poroelasticity [24, 4]. That both the stress and pressure propagate diffusively can be understood from the low-Reynolds number force balance in a gel which implies that force balance implies that pressure gradients are balanced by stress gradients. At relatively early times, when the poroelastic diffusion length $\ell_{p}$ is smaller than the domain size $\lambda$, the pressure gradient at the domain boundary can be estimated as $\nabla p\sim\delta p/\ell_{p}$; this approximation captures the transient buildup of the interfacial gradient before it reaches its long-time, saturated value. As defined by poroelastic diffusion [26]:
@@
where $D_{p}\sim k_{0}\,G/\eta_{ps}$ is the poroelastic diffusion constant, with $k_{0}$ the permeability of the gel, $G$ its elastic modulus, and $\eta_{ps}$ the viscosity associated with the viscoelastic response of the coupled polymer–solvent system activated by solvent flow at intermediate times. This description applies only until $\ell_{p}$ reaches the domain size $\lambda(t)$. Beyond this point, the pressure field has equilibrated across the entire solvent-rich domain, and extending the gradient over larger length scales does not increase the capillary driving force acting on that domain. Mathematically, this is because the finite domain size, together with the Laplace pressure boundary condition at the domain interface, fix both the magnitude and the spatial extent of the pressure gradients. Those are substantial only for distances from the domain smaller or equal to $\lambda$. As shown in more detail in Appendix C, beyond this point, the solution of the poroelastic diffusion equation has evolved so that the pressure gradient evaluated at the domain interface no longer changes with time, and further spreading of the pressure field to larger distances does not increase the capillary driving force acting on that domain.
@@
We note that most treatments of poroelasticity take the crosslinked gel to have an elastic (as opposed to viscoelastic) stress response [26, 12]. This indeed applies for times greater than the intrinsic viscoelastic crossover time of the gel, $\tau_{\mathrm{el}}$, which separates the early-time stress response, when polymer motions can relax an imposed stress, from the long-time elastic response [11, 22]. For $t\gg\tau_{\mathrm{el}}$, the shorter-time polymeric degrees of freedom have equilibrated, the crosslinked polymers behave elastically, and dissipation associated with stress redistribution is controlled primarily by solvent flow relative to a static polymer network. In this regime, the viscosity entering stress diffusion is well approximated by the Darcy viscosity, so that $\eta_{ps}\approx\eta_{D}$.
@@
During coarsening, however, the gel is responding dynamically to stresses imposed by solvent being ejected from shrinking domains and accumulated in growing domains. These stresses act as localized perturbations to a gel that is otherwise close to its equilibrium swollen state. Because these perturbations are time dependent and spatially localized, they do not probe the static, long-time elastic response of the network, but instead drive transient polymer stress relaxation. Permanent crosslinking suppresses macroscopic flow but does not eliminate internal polymer dissipation; even dry covalently crosslinked networks exhibit viscoelastic stress relaxation at intermediate times due to strand-level and non-affine relaxation modes ($\sim$ a few hundred s), before reaching an elastic plateau [25]. For this reason, we envision the gel responding viscoelastically on the timescale of coarsening, before elastic arrest intervenes and halts the dynamics at long times (on the scale of hours). The viscoelastic response of the polymer, which applies even to permanently crosslinked gels, arises from the time-dependent, local strains in the gel due, for example, from the solvent being injected by a shrinking domain. These strains generate time-dependent polymer stresses that excite internal viscoelastic relaxation modes associated with frictional rearrangements of polymer segments between crosslinks. The local and transient polymer rearrangements contribute to stress relaxation, and the relevant dissipation expressed by the effective stress-relaxation viscosity $\eta_{ps}$ reflects both solvent–polymer friction and the intrinsic viscoelastic dynamics of the polymer-rich phase [22, 18].
@@
Accordingly, we take the viscosity entering poroelastic diffusion during coarsening to be a poroelastic viscosity $\eta_{ps}$ that is distinct from the Darcy viscosity $\eta_{D}$. The Darcy viscosity characterizes solvent flow in response to pressure gradients in a static network, whereas $\eta_{ps}$ governs the coupled relaxation of pressure and polymer stress when the network itself is dynamically responding. This distinction must be made even in a permanently crosslinked gel. There too solvent transport generates strains on timescales comparable to polymer stress relaxation. In general, therefore, $\eta_{ps}\neq\eta_{D}$ for $t\ll\tau_{\mathrm{el}}$. We comment further on the implications of this distinction in the Results and Discussion.
@@
$\dot{\lambda}\sim-\frac{k_{0}\sigma}{\eta_{D}\,\lambda\,\ell_{p}},$ (3)
@@
$\lambda\sim\left(\frac{k_{0}\,\sigma}{\eta_{D}\,D_{p}^{1/2}}\right)^{1/2}t^{1/4}.$ (4)
@@
As above, shrinking and growing domains obey the same scaling by solvent conservation. In this slow-poroelastic regime, the coarsening law exhibits an explicit dependence on the gel modulus through the poroelastic diffusion constant, $D_{p}\sim G$, giving $\lambda\sim G^{-1/4}t^{1/4}$. The extent to which the prefactor $\sqrt{k_{0}\sigma/\eta_{D}}$ introduces additional modulus dependence is discussed further in the Results.

### v2

@@ Abstract
Recent experiments on crosslinked gels quenched from solvent-rich to solvent-poor conditions show solvent-rich domains embedded in a gel-rich matrix. These domains coarsen and then undergo kinetic arrest at micron scales for hours, before macroscopic drainage to equilibrium over an even longer time. Motivated by these observations, we develop a minimal model that couples capillarity-driven Darcy permeation to the viscoelastic-to-elastic crossover of the polymer network. In the viscoelastic regime, the Young–Laplace boundary traction (due to surface tension) at curved solvent–gel interfaces generates an isotropic pressure gradient in the solvent pores of the gel that drives solvent flow and coarsening. In the elastic regime, the same interfacial traction is balanced by a long-ranged deviatoric (non-isotropic) stress field, and the elastic equilibrium solution yields a spatially uniform pore pressure outside a domain, eliminating the Darcy driving force and arresting coarsening. Using the kinetic criterion $t(\lambda_{\rm arrest})\sim\tau_{\rm el}$, we predict stiffness-dependent growth laws and arrest lengths. For melt-like, polymer-rich gels, $\lambda(t)\sim G^{-1/2}t^{1/4}$ and $\lambda_{\rm arrest}\sim G^{-1/2}$. For low polymer fractions where the mesh size controls transport, $\lambda(t)\sim G^{-1/3}t^{1/3}$ and $\lambda_{\rm arrest}\sim G^{-1/3}$. The predicted $G^{-1/2}$ arrest scaling for melt-like gels is consistent with measurements.
@@ I Introduction
Recent experiments [9] on phase separation of PDMS cross-linked polymers swollen by an organic solvent (a polymer gel system) have shown results that are strikingly different from our simple understanding of phase separation in simpler, two-component systems. The cross-linked polymer PDMS gel is initially prepared at a higher temperature with a certain fraction of solvent in equilibrium. The temperature is then lowered, and the equilibrium solvent fraction within the gel decreases. Once the gel comes to true equilibrium — which only occurs after many hours — the gel and the solvent exhibit macroscopic phase separation.
At its lower temperature state, the gel initially contains excess solvent due to its initial high-temperature preparation. After an initial period of spinodal decomposition and nucleation, mesoscale domains of solvent-rich regions form in the gel, surrounded by gel that is polymer-rich and has less solvent than its initial, high temperature state. The domain structure coarsens (with small domains shrinking and large ones growing), and this process is reminiscent of Lifshitz-Slyozov-Wagner (LSW) [15, 26, 3], coarsening in the phase separation of two-component systems. Strikingly, in these experiments the coarsening is observed to undergo long-lived kinetic arrest at submicron or micron length scales, depending on the rigidity of the cross-linked gel. This mesoscale structure remains stable for several hours until, after an even longer time, all excess solvent eventually drains from the solvent-rich domains into a reservoir. The gel then becomes homogeneous at its low-temperature state with a lower solvent fraction than its prepared state.
@@
A paper with theory and model calculations of phase separation and coarsening in a viscoelastic gel by Curk and Luijten [5] extended LSW-type Ostwald ripening driven by capillarity (interfacial tension $\gamma$) by incorporating elastic stresses and their viscoelastic relaxation in the surrounding matrix (see Appendix A. Their approach is based on diffusion-controlled evaporation–condensation through a viscoelastic solid, modified by elastic and viscoelastic effects. They do not include poroelastic (Darcy) solvent flow through a deformable, porous network, which provides the dominant dissipative mechanism for relaxing capillary pressure in our approach and leads to the elastic arrest prediction of scaling with $1/G^{1/2}$, in agreement with the experiments [9]. A previous, numerical calculation by Onuki and Puri [17] focused on a two-dimensional crosslinked gel with (implicit) solvent-gel friction. In their numerical solutions the onset of elastic arrest occurred at a time that depended on the value of the modulus (see Appendix for further discussion). A more recent theory of phase separation in elastic networks [20], based on coupled chemical-potential gradients and elastic stress, concluded that local elasticity alone cannot arrest coarsening, and that non-local elastic interactions are required (see Appendix B. By contrast, our approach focuses explicitly on solvent flow through the gel coupled to the polymers’ poroelastic response. By treating solvent permeation in the gel (Darcy flow) and the elastic back-reaction directly—rather than incorporating elasticity solely through constitutive stress–energy relations—we demonstrate that elastic arrest of coarsening can occur even for linearly elastic gels, as explained below.
@@
Here, we present a unified scaling theory of both domain coarsening and elastic arrest in crosslinked gels which exhibit crosslink-controlled elastic behavior at long times and a viscoelastic response activated by solvent flow through the network at intermediate times. Solvent flow (through “Darcy’s law”) is driven by the pressure gradients within the solvent pores of the gel [25, 2, 13]. It occurs at time scales for which poroelastic dissipation, acting within the viscoelastic regime of the polymer network, controls the rate of coarsening. In contrast, the onset of long-time elastic response governs kinetic arrest [8, 17, 25], since as discussed below, the elastic equilibrium is that of a shear stress with no pressure gradient in the gel and hence no solvent flow or coarsening. When the characteristic coarsening time becomes comparable to the relaxation time separating the flowing, viscoelastic regime from the solid-like elastic regime of the crosslinked polymers, further capillary-driven domain evolution is arrested by the development of long-range elastic stress in the gel. At the level of the Laplace force balance at the solvent-rich domain boundary, the shear stress balances the capillary driving force; there is no pressure gradient in the solvent pores of the gel and no driving force for further solvent flow. The resulting arrested domain size is therefore not set directly by microscopic gel length scales, but by a kinetic criterion and is amplified by a magnification factor arising from the large separation between the poroelastic time governing solvent transport and the much longer time required for the onset of elastic response. During this extended viscoelastic window, solvent can be transferred over progressively larger distances, converting nanoscale gel properties into mesoscale domain sizes at arrest.
The complexity of the problem is simplified by our scaling approach focused on the dimensionally correct dependence of the coarsening and elastic arrest in terms of the gel properties, without the fine distinctions of numerical factors or dimensionless factors of order unity [6]. Thus, instead of equalities, we use the convention $\sim$ to mean “scales as”. The theory predicts that the coarsening length scales as $t^{1/4}$ for slow, poroelastic diffusion and that the arrest size of the domains scales as $1/\sqrt{G}$; the latter agrees with the experiments. Finally, we note here that while the explanations below consider the case of shrinking, solvent-rich domains, our model and its scaling predictions apply to both shrinking and growing domains, which are of course related by conservation of solvent. When we discuss solvent flowing out of shrinking solvent-rich domains, the same physics applies to solvent being drawn into growing domains.
@@ II.1 LSW coarsening
These domains coarsen [3] and the kinetics of coarsening are based upon the conservation relation that equates the number of molecules $N_{B}$ leaving (for example) a shrinking B domain (and eventually being drawn into a growing B domain), to their flux $J_{BA}$ at the interface of the B domain with the background of A: $dN_{B}/dt=-J_{BA}S$, where the flux is evaluated locally at the domain boundary and is controlled by the driving force acting at that boundary. For scaling purposes, we can ignore the vectors, keeping in mind that the flux and velocity of the B-molecules exiting the B-rich domain are in the direction normal to the domain boundary. Here $S$ is the area of the shrinking domain which, for simplicity, we have taken to be spherical and henceforth we denote the time derivatives by a dot. $\dot{N}_{B}$ is related to the composition difference $\Delta c$ between the two phases and the change in the domain interface position and the domain area. The flux is the number of B molecules that exit the B-rich domain, per unit area per unit time and is written in terms of the B-molecule velocity as: $J_{BA}=c_{B}\ v$, where $c_{B}$ is the concentration (number density) of B molecules in the B-rich domain. For a shrinking domain, $\dot{N}_{B}<0$ implies that $J_{BA}>0$; the velocity is directed away from the domain. In scaling, we write: $\dot{N}_{B}\sim-\Delta c\,S\,\dot{\lambda}$ where $\lambda$ is the size of the $B$ domain. For a more detailed discussion of this and the following see Appendix C.
@@
Using the conservation equation, we now write a scaling relation: $\Delta c\ \dot{\lambda}\sim-D\ \gamma/(\lambda\ \ell)$ where $\Delta c$ is the concentration difference of the B molecules in the B-rich phase and the A-rich phase. $\ell$ now is the extent of the chemical potential gradient in the direction normal to the surface of the B domain, evaluated at the domain interface. LSW scaling for the domain growth is obtained by approximating $\ell$ by $\lambda$ which originates in the solution of the diffusion equation for the B molecules in the background of A, at the domain surface [3] and Appendix C. This equation in the fast-diffusion limit (applicable to molecular diffusion in a liquid relative to the coarsening times), results in an equation for the spatial profile of the B molecule concentration in the A-rich domain where $\lambda$ determines the scale of the gradient. Below, we generalize the approximation for the gradient to include diffusion.
@@
The result for a shrinking domain gives $\dot{\lambda}\sim-(D/\Delta c)\,\gamma/\lambda^{2}$. For scaling, we can take $\Delta c\sim c_{B}$ (not valid near a critical point) which gives the LSW scaling, $\lambda\sim\left(\ (D\ \gamma/c_{B})\ t\right)^{1/3}$, where we do not distinguish shrinking from growing vis a vis the scaling; the two are simply related by conservation of the B molecules. As discussed below, the equations for the chemical potential gradients can be transformed into equations for the osmotic pressure gradients of the B molecules in the A-rich phase.
II.2 Generalizing LSW to coarsening of solvent-rich domains in viscoelastic gels:
@@
In the following, the gel properties such as interfacial tension with a solvent-rich domain, elastic modulus $G$, mesh size $\xi$ those that characterize the equilibrium gel. We again begin with the conservation relation, $\dot{N}_{B}\sim-\Delta c_{s}\,S\,\dot{\lambda}=-JS$ where the flux is evaluated locally at the domain boundary and is controlled by the driving force acting at that boundary. $\Delta c_{s}$ is the difference between the solvent concentration in the solvent-rich domain and the gel. The flux is given by $\vec{J}=c_{s}\vec{v}$ where $c_{s}$ solvent concentration in the solvent-rich domain which, due to capillary pressure, releases solvent molecules into the gel with a velocity $\vec{v}$. For a more detailed derivation of this and the following, see Appendix C.
@@
However, instead of diffusing in the polymer-rich phase, the solvent flows in the deformable, viscoelastic gel considered as a porous medium. The flow velocity $\vec{v}$ is given by the Darcy relation [6, 1, 2, 23] that relates this velocity to the pressure gradient in the solvent pores of the gel: $\vec{v}=(k_{0}/\eta_{D})\nabla p$. Here $k_{0}$ is the permeability of the gel and $\eta_{D}$ is the viscosity appropriate to the solvent flow impeded by its friction with the polymer network. Because this transport is controlled by the hydrodynamic permeability of the solvent rather than by polymer rearrangements, the network topology is effectively fixed on the Darcy-flow timescale: polymer configurations do not relax rapidly enough to eliminate the stress induced by solvent flow. Physically, this reflects a separation of timescales between the relatively fast permeation of small solvent molecules through the gel and the much slower relaxation of stress carried by long-chain polymers. In the Results, we discuss specific choices for these parameters.
@@
The pressure gradient entering Darcy’s law is the local gradient at the interface of the solvent-rich domain and the gel, whose magnitude is set by how the interfacial pressure perturbation spreads into the gel. In incompressible fluid systems pressure gradients are proportional to chemical potential gradients of the solvent molecules in the gel, by the thermodynamic relation, $\nabla\mu=\nabla p/c_{B}$; see Eq. 1.1.14 of Ref. [19] and Appendix C. The chemical potential at the solvent-rich domain surface is again given by capillarity, $\delta\mu=\gamma/(\lambda c_{s})$ and the osmotic pressure jump at the interface (which acts as the mechanical pressure jump) is given by the Laplace law as $\delta p=c_{s}\delta\mu=\gamma/\lambda$. Finally, the conservation law and the Darcy relation used in the flux give the dynamical scaling relation for a shrinking domain:
$\dot{\lambda}\sim-\frac{k_{0}\ c_{s}}{\eta_{D}}\nabla p\sim-\frac{k_{0}\ \gamma}{\eta_{D}\ \lambda\ \ell_{p}}$ (1)
@@
The solvent pressure in the gel decreases from its maximal value at the boundary of the shrinking domain via stress diffusion, known as poroelasticity [23, 4]. That both the stress and pressure propagate diffusively can be understood from the low-Reynolds number force balance in a gel where pressure gradients are balanced by viscoelastic, polymer stress gradients. At relatively early times, when the poroelastic diffusion length $\ell_{p}$ is smaller than the domain size $\lambda$, the pressure gradient at the domain boundary can be estimated as $\nabla p\sim\delta p/\ell_{p}$; this approximation captures the transient buildup of the interfacial gradient before it reaches its long-time, saturated value. As defined by poroelastic diffusion [25]:
@@
where $D_{p}\sim k_{0}\,G/\eta_{ps}$ is the poroelastic diffusion constant, with $k_{0}$ the permeability of the gel, $G$ its elastic modulus, and $\eta_{ps}$ the viscosity associated with the viscoelastic response of the coupled polymer–solvent system activated by solvent flow at intermediate times. This description applies only until $\ell_{p}$ reaches the domain size $\lambda(t)$. As shown in Appendix C, this is because the finite domain size, together with the Laplace pressure boundary condition at the domain interface, fix both the magnitude and the spatial extent of the pressure gradients. Those are substantial only for distances from the domain smaller or equal to $\lambda$. Beyond this point, the solution of the poroelastic diffusion equation has evolved so that the pressure gradient evaluated at the domain interface no longer changes with time.
@@
We note that most treatments of poroelasticity take the crosslinked gel to have an elastic (as opposed to viscoelastic) stress response [25, 12]. This indeed applies for times greater than the intrinsic viscoelastic crossover time of the gel, $\tau_{\mathrm{el}}$, which separates the early-time viscoelastic stress response, from the long-time elastic response [11, 22]. For $t\gg\tau_{\mathrm{el}}$, the shorter-time polymeric degrees of freedom have equilibrated, the crosslinked polymers behave elastically, and dissipation associated with stress redistribution is controlled primarily by solvent flow relative to a static polymer network. In this regime, the viscosity entering stress diffusion is well approximated by the Darcy viscosity, so that $\eta_{ps}\approx\eta_{D}$.
@@
During coarsening, however, the gel is responding dynamically to stresses imposed by solvent being released from shrinking domains and accumulated in growing domains. These stresses act as localized perturbations to a gel that is otherwise close to its equilibrium swollen state. Because these perturbations are time dependent and spatially localized, they do not probe the static, long-time elastic response of the network, but instead drive transient polymer stress relaxation. Permanent crosslinking suppresses macroscopic polymeric flow but does not eliminate internal polymer dissipation; even dry covalently crosslinked networks exhibit viscoelastic stress relaxation at intermediate times due to polymeric (individual chain and local, transient cooperative) relaxation modes , before reaching an elastic plateau [24]. For this reason, we envision the gel responding viscoelastically on the timescale of coarsening, before elastic arrest intervenes and halts the dynamics at long times (on the scale of hours). The viscoelastic response of the polymer, which applies even to permanently crosslinked gels, arises from the time-dependent, local strains in the gel due, for example, from the solvent being injected by a shrinking domain. These strains generate time-dependent polymer stresses that excite internal viscoelastic relaxation modes associated with frictional rearrangements of polymer segments between crosslinks. The local and transient polymer rearrangements contribute to stress relaxation, and the relevant dissipation expressed by the effective stress-relaxation viscosity $\eta_{ps}$ reflects both solvent–polymer friction and the intrinsic viscoelastic dynamics of the polymer-rich phase [22, 18].
@@
Accordingly, we take the viscosity entering poroelastic diffusion during coarsening to be a poroelastic viscosity $\eta_{ps}$ that is distinct from the Darcy viscosity $\eta_{D}$ which characterizes solvent flow in response to pressure gradients in a static network. In contrast, $\eta_{ps}$ governs the coupled relaxation of pressure and polymer stress when the network itself is dynamically responding. This distinction must be made even in a permanently crosslinked gel. There too, solvent transport generates strains on timescales comparable to polymer stress relaxation. In general, therefore, $\eta_{ps}\neq\eta_{D}$ for $t\ll\tau_{\mathrm{el}}$. We comment further on the implications of this distinction in the Results and Discussion.
@@
$\dot{\lambda}\sim-\frac{k_{0}\gamma}{\eta_{D}\,\lambda\,\ell_{p}},$ (3)
@@
$\lambda\sim\left(\frac{k_{0}\,\gamma}{\eta_{D}\,D_{p}^{1/2}}\right)^{1/2}t^{1/4}.$ (4)
@@
As above, shrinking and growing domains obey the same scaling by solvent conservation. In this slow-poroelastic regime, the coarsening law exhibits an explicit dependence on the gel modulus through the poroelastic diffusion constant, $D_{p}\sim G$, giving $\lambda\sim G^{-1/4}t^{1/4}$. The extent to which the prefactor $\sqrt{k_{0}\gamma/\eta_{D}}$ introduces additional modulus dependence is discussed further in the Results.